OpenStudy (anonymous):

Jesse played two days of golf. On the second day, he got a score of 6 below par, or –6. His total score for the two days was 0 above par, or 0. Define a variable. Then write and solve an equation to find the score Jesse got on the first day. Show your work.

OpenStudy (anonymous):

I have no idea wat this eans

OpenStudy (anonymous):

*means

OpenStudy (anonymous):

x = score on first day y = score on second day x + y = 0 y = -6 Now substitute -6 in for y in first equation x + y = 0 x + (-6) = 0 x - 6 = 0 x - 6 + 6 = 0 + 6 x = 6 check... x + y = 0 6 + (-6) = 0 6 - 6 = 0 0 = 0 (correct) :)